If a exterior angle of a regular polygon is 144°, then what is the number of its sides?

Found 2 solutions by Alan3354, ikleyn:
Answer by Alan3354(69443)

If a exterior angle of a regular polygon is 144°, then what is the number of its sides?
--------------
N = 360/Ext angle = 360/144 = 2.5 sides.
---------
If it's the interior angle = 144º, the exterior is 36º.
360/36 = 10 sides

Answer by ikleyn(52781)

.
An exterior angle of a regular polygon CAN NOT be 144°.

In other words, such a regular polygon DOES NOT EXIST.